# SMMILE: An Expert-Driven Benchmark for Multimodal Medical In-Context Learning

### v1

@@ 4.2 Analyzing Example Order
Figure 5: We analyze the effect of example order on MLLM performance. Here, we report performance across nine MLLMs (ordered by model size) in the open-ended setting with LLM-as-a-Judge evaluation.
5 Related Work
In-Context Learning: In-context learning (ICL) was popularized by brown2020language in their paper introducing GPT-3, demonstrating that LLMs can learn to solve tasks at inference time by merely conditioning on a few labeled examples in the input prompt without any gradient-based fine-tuning. This paradigm shift has enabled models to generalize to new tasks at inference time simply from natural language instructions and exemplars alone. The extension of ICL to the vision-language domain was pioneered by Flamingo, a powerful model trained on interleaved sequences of images and text (alayrac2022flamingo, ). Flamingo showcased strong few-shot performance on a wide range of visual question answering (VQA) and image captioning tasks by learning entirely from prompts composed of image-text pairs, thereby introducing the first stepping stone towards multimodal ICL.
MLLMs in Medicine:
Inspired by general-purpose MLLMs like Flamingo (alayrac2022flamingo, ) and Llava (liu2023visual, ), recent works have proposed medical MLLMs capable of handling tasks such as radiology report generation, visual question answering, and medical diagnosis. This includes works like Med-PaLM M (tu2024towards, ), Med-Flamingo (moor2023med, ), Llava-Med (li2023llava, ), ChexAgent (chen2024chexagent, ), and BiomedGPT (zhang2024generalist, ). The ability of these models to perform multimodal in-context learning has not been well studied due to a lack of available benchmarks.
Benchmarking Multimodal ICL: Evaluating the ability of MLLMs to effectively learn from multimodal in-context examples at inference time is challenging. In the general domain, several works have presented approaches for evaluating the ICL capabilities of MLLMs (zeng2024mllmsperformtexttoimageincontext, ; Baldassini_2024_CVPR, ; chen2024multimodallargelanguagemodels, ); in particular, zong2024vl recently introduced VL-ICL Bench. The domain of medicine serves as an optimal application domain for multimodal ICL due to the presence of highly-specialized concepts and complex imagery as well as the potential for real-world clinical impact. However, to the best of our knowledge, no benchmarks have been previously developed to evaluate multimodal ICL capabilities in the medical domain. Our benchmark SMMILE is designed to bridge this research gap. Prior works in the medical domain evaluated few-shot visual question answering or radiology report generation with benchmarks such as VQA-RAD (lau2018dataset, ), PathVQA (he2020pathvqa, ), SLAKE (liu2021slake, ), or MIMIC-CXR (johnson2019mimic, ); typically, few-shot examplars in these settings are chosen in an automated fashion via random selection (moor2023med, ; yan-etal-2023-style, ). In contrast, in-context examples included in SMMILE are carefully curated by experts in order to serve as relevant task demonstrations that support the learning of the task at hand.
@@ 6 Discussion
Key findings. In this work, we introduced SMMILE, a multimodal medical in-context learning benchmark designed in collaboration with a team of international clinical experts. Even the best-performing models, such as GPT-4o on SMMILE and Qwen2.5-VL-72B on SMMILE++, are only capable of answering approximately half of the questions accurately. Applying ICL results in substantial performance boosts for only a few models. Our results demonstrate a substantial gap between current MLLMs and the generalizability required for clinical use.
@@
Limitations and future work. We note several key directions for future work. (1) Scale. Crowdsourcing or synthetic augmentation could help expand coverage across the thirteen data modalities included in SMMILE. (2) Modalities. Time-series signals, volumetric scans, genomics, and structured EHR fields are not currently included. (3) Expert diversity. Current contributors may not capture all specialties or practice settings; future work can look at recruiting a larger diversity of expert contributors. (4) Task scope. The benchmark centers on diagnosis; extensions to treatment planning, prognosis, and longitudinal reasoning would help cover other aspects of clinical workflows. Nonetheless, SMMILE exposes concrete weaknesses in today’s MLLMs when applied to medical scenarios and supplies the community with a rigorous, extensible framework for further research.
@@ Acknowledgments
Appendix A Dataset Curation
In this section, we provide extended details on our four-step expert-guided data curation procedure.
A.1 Step 1: Instructions for Clinical Experts
@@
Were excited that you are participating in this research project to create a medical visual question-answering (VQA) benchmark for multimodal AI models! We focus on challenging tasks for which we provide a model with few multimodal context examples, to be followed by a final problem (see below in the visual examples) which on its own is not easy to solve for existing vision-language models (i.e., those of us with access can check with GPT4-V).
Topics: The problems can range across all medical specialties, including radiology images, photographs, pathology slices, ophthalmology imaging etc. Try to focus mostly on 2D images (e.g., slice of CT, Chest X-ray, Photograph etc.), but links to further modalities are welcome (audio, video, sequencing etc.) - as long as they can be referred to via URL.
Data: Do NOT upload any media (e.g., images, videos, audio). Instead, please provide a URL (link) to a publicly available media resource. Do not display any identifiable patient information.
Guidelines: Try to follow a consistent answer format within a given problem - if the problem allows for it. Most importantly, answers must follow a consistent format: "Epidural hematoma, left.", "Subdural hematoma, right." etc. Two in-context examples minimum - 10 maximum.
A.2 Step 2: Homepage Interface
@@
A.3 Step 3: Problem Creation
@@
A.4 Step 4: Final Submission
@@
Figure 8: After the expert clicks "Submit", they are presented with an overview of their newly created problem. The expert can then validate the problem or return to the previous screen for additional edits.
Appendix B Descriptive Statistics for SMMILE++
@@
Figure 9: Dataset characteristics. (A–D) Distribution of four key categorical annotations across the unique problems: (A) answer format, (B) rarity of the clinical case based on how often clinicians would experience the medical concepts included in each problem, (C) primary cognitive process required (where reasoning classification is defined by final problem not having direct support in its in-context example set), and (D) rated difficulty for state-of-the-art LLMs based on model performance. (E–F) Horizontal barplots showing the breakdown of each problem by its main medical specialty (E) and by main image type used (F). (G) Histogram of the number of in-context examples provided per problem. (H) Overlaid histograms of the character-length distributions for questions versus answers. All panels are based on the 1038 problems included in SMMILE++.
Appendix C Additional Experimental Details
@@
LLM-as-a-Judge evaluations were performed with Llama3.3 70B accessed via the Ollama software package [^2] Ollama can be accessed at https://github.com/ollama/ollama.. The input prompt is provided below. The output is a binary value, which we multiply by 100 to achieve a final score of either 0 (incorrect) or 100 (correct) for each generated output.
A medical AI model is provided with an image and asked the question "{question}". The correct answer to this question is: "{answer}". The AI model outputs "{response}" as its response. Is the AI model correct? Please output your answer as a single digit, where 1 indicates that the AI model is correct and 0 indicates that the AI model is incorrect with respect to the correct answer. Do not provide anything other than the digit in your response.
Appendix D Extended Fine-Grained Analysis for SMMILE
@@
Appendix E Extended Fine-Grained Analysis for SMMILE++
@@
Appendix F Licensing Considerations

### v2

@@ 4.2 Analyzing Example Order
Figure 5: We analyze the effect of example order on MLLM performance. We report performance across 9 MLLMs (ordered by model size) in the open-ended setting with LLM-as-a-Judge evaluation.
@@ 5 Discussion
Key findings. In this work, we introduced SMMILE, a multimodal medical in-context learning benchmark designed in collaboration with a team of international clinical experts. Even the best-performing models, such as GPT-4o on SMMILE and Qwen2.5-VL-72B on SMMILE++, are only capable of answering approximately half of the questions accurately. Applying ICL results in substantial performance boosts for only a few models. Our results demonstrate a significant gap between current MLLMs and the generalizability required for clinical use. Limitations and future work are discussed in Appendix Section G.
@@ Acknowledgments
Appendix A Related Work
In-Context Learning: In-context learning (ICL) was popularized by [5] in their paper introducing GPT-3, demonstrating that LLMs can learn to solve tasks at inference time by merely conditioning on a few labeled examples in the input prompt without any gradient-based fine-tuning. This paradigm shift has enabled models to generalize to new tasks at inference time simply from natural language instructions and exemplars alone. The extension of ICL to the vision-language domain was pioneered by Flamingo, a powerful model trained on interleaved sequences of images and text [1]. Flamingo showcased strong few-shot performance on a wide range of visual question answering (VQA) and image captioning tasks by learning entirely from prompts composed of image-text pairs, thereby introducing the first stepping stone towards multimodal ICL.
MLLMs in Medicine:
Inspired by general-purpose MLLMs like Flamingo [1] and Llava [24], recent works have proposed medical MLLMs capable of handling tasks such as radiology report generation, visual question answering, and medical diagnosis. This includes works like Med-PaLM M [28], Med-Flamingo [25], Llava-Med [20], ChexAgent [7], and BiomedGPT [35]. The ability of these models to perform multimodal in-context learning has not been well studied due to a lack of available benchmarks.
Benchmarking Multimodal ICL: Evaluating the ability of MLLMs to effectively learn from multimodal in-context examples at inference time is challenging. In the general domain, several works have presented approaches for evaluating the ICL capabilities of MLLMs [34, 4, 6]; in particular, [37] recently introduced VL-ICL Bench. The domain of medicine serves as an optimal application domain for multimodal ICL due to the presence of highly-specialized concepts and complex imagery as well as the potential for real-world clinical impact. However, to the best of our knowledge, no benchmarks have been previously developed to evaluate multimodal ICL capabilities in the medical domain. Our benchmark SMMILE is designed to bridge this research gap. Prior works in the medical domain evaluated few-shot visual question answering or radiology report generation with benchmarks such as VQA-RAD [14], PathVQA [11], SLAKE [21], or MIMIC-CXR [12]; typically, few-shot examplars in these settings are chosen in an automated fashion via random selection [25, 32]. In contrast, in-context examples included in SMMILE are carefully curated by experts in order to serve as relevant task demonstrations that support the learning of the task at hand.
Appendix B Dataset Curation
In this section, we provide extended details on our four-step expert-guided data curation procedure. The front-end of our data collection platform consists of a single‑page React client that collects each panel’s metadata (question, answer, public image URL, specialty, author, order). Once the contributor finishes a problem, the client posts the structured annotations to the back‑end. The back-end converts these annotations to an parquet file and uploads the shard to a version‑controlled HuggingFace Hub dataset.
B.1 Step 1: Instructions for Clinical Experts
@@
Instructions for Clinical Experts We’re excited that you are participating in this research project to create a medical visual question-answering (VQA) benchmark for multimodal AI models! We focus on challenging tasks for which we provide a model with few multimodal context examples, to be followed by a final problem (see below in the visual examples) which on its own is not easy to solve for existing vision-language models (i.e., those of us with access can check with GPT4-V). Topics: The problems can range across all medical specialties, including radiology images, photographs, pathology slices, ophthalmology imaging etc. Try to focus mostly on 2D images (e.g., slice of CT, Chest X-ray, Photograph etc.), but links to further modalities are welcome (audio, video, sequencing etc.) - as long as they can be referred to via URL. Data: Do NOT upload any media (e.g., images, videos, audio). Instead, please provide a URL (link) to a publicly available media resource. Do not display any identifiable patient information. Guidelines: Try to follow a consistent answer format within a given problem - if the problem allows for it. Most importantly, answers must follow a consistent format: "Epidural hematoma, left.", "Subdural hematoma, right." etc. Two in-context examples minimum - 10 maximum.
B.2 Step 2: Homepage Interface
@@
B.3 Step 3: Problem Creation
@@
B.4 Step 4: Final Submission
@@
Figure 8: After the expert clicks "Submit", they are presented with an overview of their newly created problem. The expert can then validate the problem or return to the previous screen for more edits.
Appendix C Descriptive Statistics for SMMILE++
@@
Figure 9: Dataset characteristics. (A–D) Distribution of four key categorical annotations across the unique problems: (A) answer format, (B) rarity of the clinical case based on how often clinicians would experience the medical concepts included in each problem, (C) primary cognitive process required (where reasoning classification is defined by final problem not having direct support in its in-context example set), and (D) rated difficulty. (E–F) Horizontal barplots showing the breakdown of each problem by its main medical specialty (E) and by main image type used (F). (G) Histogram of the number of in-context examples provided per problem. (H) Overlaid histograms of the character-length distributions for questions versus answers. All panels are based on the 1038 problems included in SMMILE++.
Appendix D Additional Experimental Details
D.1 Computational Requirements
@@
D.2 LLM-as-a-Judge Implementation Details
The LLM-as-a-Judge evaluations capture semantic correctness beyond exact string matching, making it particularly valuable for medical reasoning tasks where multiple phrasings might convey the same diagnosis or finding. LLM-as-a-Judge evaluations were performed with Llama3.3 70B, accessed via the Ollama software package [^2] Ollama can be accessed at https://github.com/ollama/ollama.. The input prompt is provided below. The output is a binary value, which we multiply by 100 to achieve a final score of either 0 (incorrect) or 100 (correct) for each generated output.
LLM-as-a-Judge prompt A medical AI model is provided with an image and asked the question "question". The correct answer to this question is: "answer". The AI model outputs "response" as its response. Is the AI model correct? Please output your answer as a single digit, where 1 indicates that the AI model is correct and 0 indicates that the AI model is incorrect with respect to the correct answer. Do not provide anything other than the digit in your response.
We opted to run LLM-as-a-Judge evaluations with Llama3.3 70B because (1) Llama3.3 has been shown in prior work [9] to demonstrate strong performance on textual analysis tasks, and (2) Llama3.3 is open-source, generates reproducible results, and does not require payment, ensuring that our benchmark can be useful even in resource-constrained settings. Using a stronger LLM such as GPT-4o results in largely similar results to using Llama3.3, with high interrater agreement observed between the two models (Cohen’s kappa = 0.943 across results from six MLLMs in the open-ended setting).
D.3 Analysis of Prompting Approach
For all MLLMs evaluated in this work, we use a standard input prompt consisting of a system message, in-context examples, and the query image and question. In this section, we provide an analysis of input prompt structure on performance; specifically, we compare our standard prompting approach with Chain-of-Thought (CoT) prompting [29]. CoT prompting operates as follows: for each problem in the SMMILE benchmark, we present a system message and the multimodal in-context examples to the MLLM, followed by a query consisting of an image, a question, and an instruction of the form, "First, explain your reasoning step-by-step by referring to the provided image. Then, answer the question." Results on the SMMILE benchmark (open-ended ICL setting) are summarized in Table 4.
Table 4: Here, we analyze the effects of prompt structure on SMMILE benchmark performance (open-ended ICL setting) across a sample of 5 MLLMs. We consider two options for prompt structure: standard prompting and chain-of-thought (CoT) prompting.
Model Standard Prompting CoT Prompting
LLaVA-OneVision-0.5B $\textbf{21.63}\pm\scriptstyle{4.00}$ $7.23\pm\scriptstyle{2.53}$
Qwen2.5-VL-3B $\textbf{33.58}\pm\scriptstyle{4.09}$ $27.63\pm\scriptstyle{4.91}$
LLaVA-Med-7B $10.19\pm\scriptstyle{3.06}$ $\textbf{12.45}\pm\scriptstyle{3.14}$
Qwen2.5-VL-7B $29.58\pm\scriptstyle{4.63}$ $\textbf{29.95}\pm\scriptstyle{4.75}$
LLaVA-v1.5-13B $\textbf{20.91}\pm\scriptstyle{3.49}$ $15.75\pm\scriptstyle{3.44}$
Across the evaluated models, we see that performance improvements afforded by CoT are minor, and in fact, multiple models exhibit degraded performance when using CoT prompting. In particular, we observe substantial drops in performance for LLaVA-OneVision-0.5B and LLaVA-v1.5-13B. Further analysis demonstrates that both models exhibit high rates of malformed outputs (e.g. outputs such as "ooooooooooo…" or "( and ( ( (and (…"), suggesting that these models are unable to effectively respond to the prompt. Additionally, using CoT prompting results in substantial increases in inference time, particularly for the Qwen model family. As a result, we utilize the standard prompting approach for all evaluations in this work.
Appendix E Extended Fine-Grained Analysis for SMMILE
@@
Appendix F Extended Fine-Grained Analysis for SMMILE++
@@
Appendix G Limitations and Future Work
We note several key directions for future work:
  1. 1.
    Scale. Crowdsourcing or synthetic augmentation could help expand coverage across the thirteen data modalities included in SMMILE.
  2. 2.
    Modalities. Time-series signals, volumetric scans, genomics, and structured EHR fields are not currently included.
  3. 3.
    Expert diversity. Current contributors may not capture all specialties or practice settings; future work can look at recruiting a larger diversity of expert contributors.
  4. 4.
    Task scope. The benchmark centers on diagnosis; extensions to treatment planning, prognosis, and longitudinal reasoning would help cover other aspects of clinical workflows.
Nonetheless, SMMILE exposes concrete weaknesses in today’s MLLMs when applied to medical scenarios and supplies the community with a rigorous, extensible framework for further research.
Appendix H Licensing Considerations